OpenStudy (anonymous):

Does the equation represent a direct variation? -4x + 3y = 0 Yes, 4 Yes, 3/4 Yes, 4/3 Yes, - 4/3 no

OpenStudy (stamp):

\[-4x+3y=0\] \[3y=4x\] \[y=4x/3\] Yes, 4/3. For every y increase of 4, x will increase by 3. Example: \[y=4x/3\] \[f(x)=4x/3\] \[f(0)=4(0)/3=0/3=0\] f(0+3)=f(3)=4(3)/3=12/3=4 (0,0) to (3,4). This function as linear, or line equation. If you show it on graph paper it makes a straight line.